# Non-commutative rings with infinitely many maximal subrings

### v1

@@ 4. Rings integral over their centers
$(1)$ Let $J$ be the set of all $i\in I$ such that $T_{i}$ is not a quasi duo ring. Now we have two cases. First assume that $J$ is infinite. For each $j\in J$, let $M_{i}$ be a maximal one-sided ideal of $T_{j}$ which is not an ideal of $T_{j}$, and $N_{j}=\prod_{i\in I}A_{i}$, where $A_{i}=T_{i}$ for $i\neq j$ and $A_{j}=M_{j}$. Clearly, each $N_{j}$ is a maximal one-sided ideal of $T$ which is not an ideal of $T$. Hence $T$ has infinitely many maximal one-sided ideal which are not ideals of $T$. Thus $T$ has infinitely many maximal subrings by $(1)$ of Proposition 3.10. Hence assume that $J$ is finite. Without lose of the generality we may assume that $J=\emptyset$, i.e., each $T_{i}$ is a quasi duo ring and each $T_{i}$ has only finitely many maximal subrings. Therefore by Proposition 4.1, for each $i\in I$, the ring $T_{i}/J(T_{i})$ is a commutative ring. Take $J:=\prod_{i\in I}J(T_{i})$, then clearly $T/J\cong\prod_{i\in I}T/J(T_{i})$. Thus $T/J$ is a product of an infinite family of commutative rings, and therefore by [9, Remark 3.18] (or see [10, Theorem 3.9]), $T/J$ has infinitely many maximal subrings and so does $T$. Hence $(1)$ holds. For $(2)$, it suffices to show that each $T_{i}$ is integral over its center and use $(1)$. Clearly, $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$. Let $k\in I$ and $t\in T_{k}$, and put $x=(x_{i})_{i\in I}$ where $x_{i}=0$ for $i\neq k$ and $x_{k}=t$. By our assumption, $x$ is integral over the center of $T$. Hence there exist $n\in\mathbb{N}$ and $c_{0},c_{1},\ldots,c_{n-1}$ in $C_{T}(T)$ such that $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$. Since $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$, we have $c_{j}=(c_{ji})_{i\in I}$, for $0\leq j\leq n-1$, where $c_{ji}\in C_{T_{i}}(T_{i})$. Therefore by calculating the $k$-th component from $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$, we immediately conclude that $t^{n}+c_{n-1k}t^{n-1}+\cdots+c_{1k}t+c_{0k}=0$. Hence $t$ is integral over $C_{T_{i}}(T_{i})$ and we are done by $(1)$. ∎
@@ 5. Direct product and matrix rings
In this section we study the number of maximal subrings in matrix rings and (certain) direct products of rings. In [10, Theorem 3.9], the authors exactly determined when a direct product of a family of commutative rings has only finitely many maximal subrings. In particular, if $T$ is an infinite direct product of a family of commutative rings $\{R_{i}\}_{i\in I}$, then $|RgMax(T)|\geq 2^{|I|}$. As we see in the final result of the previous section, if $T$ is a direct product of an infinite family $\{T_{i}\}_{i\in I}$, and either each $T_{i}$ is integral over its center, or $T$ is integral over its center, then $T$ has infinitely many maximal subrings. This immediately implies that whenever for each $i$, $T_{i}=\mathbb{M}_{n_{i}}(E_{i})$, where $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, then $T$ has infinitely many maximal subrings (note, for each $i$, $\mathbb{M}_{n_{i}}(E_{i})$ is finite dimensional over its center and therefore is integral over its center). In the next proposition we exactly determine when $T$ has only finitely many maximal subrings.
@@
    For each $i,j\in I$, where $n_{i}=n_{j}=1$, and $E_{i}$ is infinite, then $E_{i}\ncong E_{j}$.
@@
Therefore, if $T$ has only finitely many maximal subrings, then $T=E_{1}\times\cdots\times E_{r}\times\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, where $n_{i}>1$ and $r,s\geq 0$. In particular, $T$ is a semisimple ring. Moreover, if $A=E_{1}\times\cdots\times E_{r}$ and $B=\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, then each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$.
@@
First assume that $T$ has only finitely many maximal subrings. This immediately implies that each $\mathbb{M}_{n_{i}}(E_{i})$ has finitely many maximal subrings. Hence if $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings and if $n_{i}>1$, then by Theorem 3.5, $\mathbb{M}_{n_{i}}(E_{i})$ is finite and therefore $E_{i}$ is finite too. Therefore $(2)$ and $(4)$ hold. Although the proof of $(1)$ is obvious by the comment preceding the proposition, but we give another proof too. To see this, let $J:=\{i\in I\ |\ n_{i}=1\}$, $J^{\prime}=I\setminus J$, $A=\prod_{i\in J}E_{i}$ and $B=\prod_{i\in J^{\prime}}\mathbb{M}_{n_{i}}(E_{i})$. Then clearly $T=A\times B$. Hence by our assumption $A$ and $B$ has only finitely many maximal subrings. Thus by [10, Theorem 3.9], we immediately conclude that $J$ is finite. It is clear that by $(5)$ of Proposition 3.10, $J^{\prime}$ is finite and therefore $I$ is finite. Finally note that $(3)$ is an immediate by Proposition 3.6. Conversely, assume that (1)(4)(1)-(4) hold and let $A$ and $B$ are defined as above. We prove that $T$ has only finitely many maximal subrings. To see this first note that $A$ and $B$ are homomorpically non-isomorphic rings, hence by Lemma 2.9, each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$. This immediately shows that $T$ has only finitely many maximal subrings, for $B$ is finite and by [10, Corollary 3.12 and $(1)$ of Proposition 3.15], $A$ has only finitely many maximal subrings. Thus we are done. ∎
@@
Assume that $T$ has only finitely many maximal subrings. Then similar to the proof of the previous theorem we deduce that $I$ is finite and for each $i\in I$, $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings. Hence by Theorem 3.5, $D_{i}$ is finite and thus $T$ is finite. The converse is evident. ∎
@@
    for each $i\neq j$ in $I$, whenever $n_{i}=n_{j}=1$ and $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$.
@@
First assume that $T$ has only finitely many maximal subrings. Thus for each $i\in I$, we conclude that $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings. Now if $n_{i}>1$, then by Theorem 3.5 we deduce that $D_{i}$ is a finite field. Hence assume that $n_{i}=1$, we claim that $D_{i}$ is a field. For otherwise by Proposition 3.1, $D_{i}$ has infinitely many maximal subrings, which is impossible. Thus for each $i\in I$, we conclude that $D_{i}$ is a field and hence we are done by Proposition 5.1. ∎
In [10, Corollary 3.10], the authors proved that if $R$ is a commutative ring, then $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring (i.e., has only finitely many maximal ideals) with only finitely many maximal subrings (in particular, $R$ is integral over $\mathbb{Z}_{n}$ for some $n>1$) and for each maximal ideal $M$ of $R$, the field $R/M$ is finite. For non-commutative rings we have the following result.
@@
    $Max(R)$, $Max_{l}(R)$ and $Max_{r}(R)$ are finite sets. In particular, $Max(T)$, $Max_{l}(T)$ and $Max_{r}(T)$ are finite sets.
@@
    For each $M\in Max(R)$, $R/M$ is finite. In particular, for each maximal ideal $Q$ of $T$, the ring $T/Q$ is finite.
@@
    For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite). In particular, for each maximal one-sided ideal $N$ of $T$, $T/M$ is finite (i.e., each simple left/right $T$-module is finite).
@@
First note that for each maximal ideal $M$ of $R$, $T/(M\times M)\cong R/M\times R/M$ has only finitely many maximal subrings. Thus by Proposition 3.6, $R/M$ is finite. Hence the first part of $(2)$ holds. Also note that for each $M\in Max(R)$, $T$ has a maximal subring $R_{M}$ which contains $M\times M$ (by the comments preceding Theorem 2.7) and clearly whenever $M\neq N\in Max(R)$, then $R_{M}\neq R_{N}$, for $(M\times M)+(N\times N)=T$. Hence we conclude that $Max(R)$ is finite. Since $T$ has only finitely many maximal subrings, we infer that $R$ has only finitely many maximal subring. This immediately implies that $Max_{l}(R)\setminus Max(R)$ is finite, by $(1)$ of Proposition 3.10. Therefore $Max_{l}(R)$ is finite, for $Max(R)$ is finite by the first part of the proof. Similarly $Max_{r}(R)$ is finite. For the second part of $(1)$, note that maximal (left/right) ideals of $T$ are of the forms $M\times R$ or $R\times M$, where $M$ is a maximal (left/right) ideal of $R$. Hence the conclusions are immediate by the first part of $(1)$. By the latter comment and the first part of $(2)$, we deduce the second part of $(2)$. For $(3)$, note that since $R$ has only finitely many maximal left ideals, we conclude that $R/J(R)$ is a semisimple ring. Hence by $(2)$, we immediately conclude that $R/J(R)$ is finite. The second part of $(3)$ is evident, for $J(T)=J(R)\times J(R)$. $(4)$ is an immediate consequences of $(3)$, for $R/J(R)$ (resp. $T/J(T)$) is a finite ring. ∎
By the previous result we immediately conclude that if $R$ is a ring and $T=R\times R$ has only finitely many maximal subrings, then there exists a finite number of non-isomorphic finite simple left $R$-module $S_{1},\ldots,S_{n}$ such that each simple left $R$-module is isomorphic to one of the $S_{i}$.
@@
    $Max(R)$ and therefore $Max(T)$ are finite. Moreover, $Max_{l}(R)$ and $Max_{r}(R)$ are finite.
@@
    $R/J(R)$ and therefore $T/J(T)$ are finite semisimple ring. In particular, $Max_{l}(T)$ and $Max_{r}(T)$ are finite.
@@
    For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite). In particular, for each maximal one-sided ideal $N$ of $T$, $T/M$ is finite (i.e., each simple left/right $T$-module is finite).
@@
Clearly for each maximal subring $S$ of $R$, $\mathbb{M}_{n}(S)$ is a maximal subring of $T$, thus we deduce that $R$ has only finitely many maximal subrings. Hence $(1)$ holds. For $(2)$, note that for each maximal ideal $M$ of $R$, $A_{M}=\mathbb{M}_{n}(M)$ is a maximal ideal of $T$ and $T/A_{M}\cong\mathbb{M}_{n}(R/M)$. By the assumption that $T$ has only finitely many maximal subrings, we deduce that $T/A_{M}$ has only finitely many maximal subrings too. Hence by Theorem 3.5, we deduce that $R/M$ is finite (note, $n>1$ and therefore $T/A_{M}$ is not a division ring). The second part of $(2)$ is obvious by the first part of $(2)$, for maximal ideals of $T$ is of the form $A_{M}$, for some maximal ideal $M$ of $R$. For $(3)$, by the notation of $(2)$, since $n>1$, by $(4)$ of Theorem 2.7, we conclude that $\mathbb{M}_{n}(R/M)$ has a maximal subring. Hence, $T$ has a maximal subring which contains $A_{M}$, say $T_{M}$. In particular, if $M\neq N\in Max(R)$, then $T_{M}\neq T_{N}$, for $A_{M}+A_{N}=T$. Thus we deduce that $Max(R)$ is finite. Therefore $Max(T)$ is finite too. For the second part of $(3)$, note that by $(1)$, $R$ has only finitely many maximal subrings, hence by $(1)$ of Proposition 3.10, we conclude that $Max_{l}(R)\setminus Max(R)$ and $Max_{r}(R)\setminus Max(R)$ are finite. Therefore $Max_{l}(R)$ and $Max_{r}(R)$ are finite, for $Max(R)$ is finite by the first part of $(3)$. For $(4)$, note that since $Max_{l}(R)$ is finite we deduce that $R/J(R)$ is a semisimple ring. Since for each maximal ideal $M$ of $R$, $R/M$ is a finite ring, we immediately conclude that $R/J(R)$ is a finite ring. Since $J(T)=\mathbb{M}_{n}(J(R))$, we deduce that $T/J(T)$ is finite semisimple too. Therefore $Max_{l}(T)$ and $Max_{r}(T)$ are finite. $(5)$ is obvious by $(4)$. ∎
Now we have the following corollaries.
@@
Let $R$ be a zero-dimensional ring with $Char(R)=0$. Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ has infinitely many maximal subrings.
@@
Assume that $\mathbb{M}_{n}(R)$ or $R\times R$ has only finitely many maximal subrings. Then by Propositions 5.4 or 5.5, we deduce that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring and therefore $Char(R/M)\neq 0$. Now note that there exists a prime ideal $P$ of $R$ such that $P\cap\mathbb{Z}=0$. Since $R$ is a zero-dimensional ring we deduce that $P$ is a maximal ideal of $R$ and therefore $R/P$ is finite. But from $\mathbb{Z}\cap P=0$, we deduce that $R/P$ contains a copy of $\mathbb{Z}$, which is absurd. ∎
@@
Let $T$ be a semilocal ring (in particular, if $T$ is a left/right Artinian ring) with only finitely many maximal subrings. Then $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$, $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$, for some $n,k\geq 0$. Moreover, whenever $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$ for $j\neq i$.
@@
Since $T/J(T)$ is a semisimple ring, we deduce that $T/J(T)\cong\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where $I$ is finite, $D_{i}$ are division rings and $n_{i}\in\mathbb{N}$. Clearly $T$ has only finitely many maximal ideals. Also note that since $T$ has only finitely many maximal subrings, then by $(1)$ of Proposition 3.10, we conclude that $Max_{l}(T)\setminus Max(T)$ is finite. Therefore $Max_{l}(T)$ is finite. Similarly $Max_{r}(R)$ is finite. Hence by Corollary 2.4, we deduce that whenever $i\in I$ and $n_{i}>1$, then $D_{i}$ is finite, i.e., $D_{i}$ is a finite field. Hence we infer that $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$ and $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$ for some $n,k\geq 0$. Finally if for some $1\leq i\neq j\leq n$, $D_{i}\cong D_{j}$, then $T$ has an ideal $I$ such that $T/I\cong D_{i}\times D_{j}\cong D_{i}\times D_{i}$. Since $T$ has only finitely many maximal subrings, we deduce that $T/I$ has only finitely many maximal subrings. Hence by Proposition 3.6, $D_{i}$ is finite. Hence we are done. ∎
We conclude this paper by the following result.

### v2

@@ 4. Rings integral over their centers
$(1)$ Let $J$ be the set of all $i\in I$ such that $T_{i}$ is not a quasi-duo ring. Now we have two cases. First assume that $J$ is infinite. Without loss of generality we may suppose that for each $j\in J$, the ring $T_{j}$ is not a left quasi-duo ring. Consequently, for each $j\in J$, let $M_{i}$ be a maximal left ideal of $T_{j}$ which is not an ideal of $T_{j}$, and $N_{j}=\prod_{i\in I}A_{i}$, where $A_{i}=T_{i}$ for $i\neq j$ and $A_{j}=M_{j}$. Clearly, each $N_{j}$ is a maximal left ideal of $T$ which is not an ideal of $T$. Hence $T$ has infinitely many maximal left ideal which are not ideals of $T$. Thus $T$ has infinitely many maximal subrings by Proposition 3.10(1)(1). Now, assume that $J$ is finite. Without loss of generality we may suppose that $J=\emptyset$, i.e., each $T_{i}$ is a quasi-duo ring with only finitely many maximal subrings. Therefore by Proposition 4.1, for each $i\in I$, the ring $T_{i}/J(T_{i})$ is a commutative ring. Take $J:=\prod_{i\in I}J(T_{i})$, then clearly $T/J\cong\prod_{i\in I}T/J(T_{i})$. Thus $T/J$ is a product of an infinite family of commutative rings. Consequently, by [9, Remark 3.18] (or see [10, Theorem 3.9]), $T/J$ has infinitely many maximal subrings and so does $T$. Hence $(1)$ holds.
$(2)$ It suffices to show that each $T_{i}$ is integral over its center and use $(1)$. Clearly, $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$. Let $k\in I$ and $t\in T_{k}$, and put $x=(x_{i})_{i\in I}$ where $x_{i}=0$ for $i\neq k$ and $x_{k}=t$. By our assumption, $x$ is integral over the center of $T$. Hence there exist $n\in\mathbb{N}$ and $c_{0},c_{1},\ldots,c_{n-1}$ in $C_{T}(T)$ such that $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$. Since $C_{T}(T)=\prod_{i\in I}C_{T_{i}}(T_{i})$, we have $c_{j}=(c_{ji})_{i\in I}$, for $0\leq j\leq n-1$, where $c_{ji}\in C_{T_{i}}(T_{i})$. Now by calculating the $k$-th component of $x^{n}+c_{n-1}x^{n-1}+\cdots+c_{1}x+c_{0}=0$, we obtain that $t^{n}+c_{n-1k}t^{n-1}+\cdots+c_{1k}t+c_{0k}=0$. It follows that $t$ is integral over $C_{T_{i}}(T_{i})$ and we are done by $(1)$. ∎
@@ 5. Direct product and matrix rings
In this section we study the number of maximal subrings in matrix rings and in certain direct products of rings. In [10, Theorem 3.9], the authors precisely determined when a direct product of a family of commutative rings has only finitely many maximal subrings. In particular, if $T$ is an infinite direct product of a family of commutative rings $\{R_{i}\}_{i\in I}$, then $|\operatorname{RgMax}(T)|\geq 2^{|I|}$. As shown in the final result of the previous section, if $T=\prod_{i\in I}T_{i}$ is a direct product of an infinite family of rings, and either each $T_{i}$ is integral over its center, or $T$ itself is integral over its center, then $T$ has infinitely many maximal subrings. This immediately implies that whenever each $T_{i}=\mathbb{M}_{n_{i}}(E_{i})$, with $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, then $T$ has infinitely many maximal subrings (note that each $\mathbb{M}_{n_{i}}(E_{i})$ is finite-dimensional over its center and therefore integral over its center). In the next proposition we give precise characterization of when such a product has only finitely many maximal subrings.
@@
    For each $i\neq j$ in $I$, if $n_{i}=n_{j}=1$, and $E_{i}$ is infinite, then $E_{i}\ncong E_{j}$, as rings.
@@
Consequently, if $T$ has only finitely many maximal subrings, then $T=E_{1}\times\cdots\times E_{r}\times\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, where $n_{i}>1$ and $r,s\geq 0$. In particular, $T$ is a semisimple ring. Moreover, if $A=E_{1}\times\cdots\times E_{r}$ and $B=\mathbb{M}_{n_{1}}(E_{r+1})\times\cdots\mathbb{M}_{n_{s}}(E_{r+s})$, then each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$.
@@
First assume that $T$ has only finitely many maximal subrings. This immediately implies that each $\mathbb{M}_{n_{i}}(E_{i})$ has finitely many maximal subrings. Hence if $n_{i}=1$, $E_{i}$ has only finitely many maximal subrings and if $n_{i}>1$, then by Theorem 3.5, $\mathbb{M}_{n_{i}}(E_{i})$ is finite and at once $E_{i}$ is finite too. Therefore $(2)$ and $(4)$ hold. Although the proof of $(1)$ is obvious by the comment preceding the proposition, but we give another proof. To see this, let $J:=\{i\in I\ |\ n_{i}=1\}$, $J^{\prime}=I\setminus J$, $A=\prod_{i\in J}E_{i}$ and $B=\prod_{i\in J^{\prime}}\mathbb{M}_{n_{i}}(E_{i})$. Then clearly $T=A\times B$. Consequently by our assumption $A$ and $B$ has only finitely many maximal subrings. Thus by [10, Theorem 3.9], we obtain that $J$ is finite. It is clear that by $(5)$ of Proposition 3.10, $J^{\prime}$ is finite and therefore $I$ is finite. Finally note that $(3)$ is an immediate consequence of Proposition 3.6. Conversely, suppose that (1)(4)(1)-(4) hold and let $A$ and $B$ are defined as above. We prove that $T$ has only finitely many maximal subrings. To see this first note that $A$ and $B$ are homomorphically non-isomorphic rings, hence by Lemma 2.9, each maximal subring of $T$ is of the form $A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$, or $A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$. This implies that $T$ has only finitely many maximal subrings, for $B$ is finite and by [10, Corollary 3.12 and of Proposition 3.15(1)(1)], AA has only finitely many maximal subrings. Thus we are done. ∎
@@
Assume that $T$ has only finitely many maximal subrings. Then analogous to the proof of the previous theorem we deduce that $I$ is finite and for each $i\in I$, $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings. Hence by Theorem 3.5, $D_{i}$ is finite and thus $T$ is finite. The converse is evident. ∎
@@
    For each $i\neq j$ in $I$, if $n_{i}=n_{j}=1$ and $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$, as rings.
@@
First assume that $T$ has only finitely many maximal subrings. Thus for each $i\in I$, we conclude that $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings. Now if $n_{i}>1$, then by Theorem 3.5 we deduce that $D_{i}$ is a finite field. Hence assume that $n_{i}=1$, we claim that $D_{i}$ is a field. Otherwise by Proposition 3.1, $D_{i}$ has infinitely many maximal subrings, which is impossible. Thus for each $i\in I$, we conclude that $D_{i}$ is a field and hence we are done by Proposition 5.1. ∎
In [10, Corollary 3.10], the authors proved that for a commutative ring $R$, the ring $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring (i.e., has only finitely many maximal ideals) with only finitely many maximal subrings (in particular, $R$ is integral over $\mathbb{Z}_{n}$ for some $n>1$) and for ever maximal ideal $M$ of $R$, the residue field $R/M$ is finite. For noncommutative rings we obtain the following generalization.
@@
    $\operatorname{Max}(R)$, $\operatorname{{Max}_{\ell}}(R)$ and $\operatorname{{Max}_{r}}(R)$ are finite sets. In particular, $\operatorname{Max}(T)$, $\operatorname{{Max}_{\ell}}(T)$ and $\operatorname{{Max}_{r}}(T)$ are finite sets.
@@
    For each $M\in\operatorname{Max}(R)$, $R/M$ is finite. In particular, for each maximal ideal $Q$ of $T$, the ring $T/Q$ is finite.
@@
    For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite). In particular, for each maximal one-sided ideal $N$ of $T$, $T/N$ is finite (i.e., each simple left/right $T$-module is finite).
@@
First note that for each maximal ideal $M$ of $R$, $T/(M\times M)\cong R/M\times R/M$ has only finitely many maximal subrings. Thus by Proposition 3.6, $R/M$ is finite. Hence the first part of $(2)$ holds. Also note that for each $M\in\operatorname{Max}(R)$, $T$ has a maximal subring $R_{M}$ which contains $M\times M$ (by the comments preceding Theorem 2.7) and clearly whenever $M\neq N\in\operatorname{Max}(R)$, then $R_{M}\neq R_{N}$, because $(M\times M)+(N\times N)=T$. Hence we conclude that $\operatorname{Max}(R)$ is finite. Since $T$ has only finitely many maximal subrings, we infer that $R$ has only finitely many maximal subring. This implies that $\operatorname{{Max}_{\ell}}(R)\setminus\operatorname{Max}(R)$ is finite, by $(1)$ of Proposition 3.10. Consequently $\operatorname{{Max}_{\ell}}(R)$ is finite, because $\operatorname{Max}(R)$ is finite by the first part of the proof. Analogously $\operatorname{{Max}_{r}}(R)$ is finite. For the second part of $(1)$, note that maximal (left/right) ideals of $T$ are of the forms $M\times R$ or $R\times M$, where $M$ is a maximal (left/right) ideal of $R$. Hence the conclusions are immediate by the first part of $(1)$. By the latter comment and the first part of $(2)$, we deduce the second part of $(2)$. For $(3)$, note that since $R$ has only finitely many maximal left ideals, we conclude that $R/J(R)$ is a semisimple ring. Hence by $(2)$, we have $R/J(R)$ is finite. The second part of $(3)$ is evident, because $J(T)=J(R)\times J(R)$. $(4)$ is an immediate consequences of $(3)$, because $R/J(R)$ and $T/J(T)$ are finite rings. ∎
From the previous result we immediately conclude that if $R$ is a ring for which $T=R\times R$ has only finitely many maximal subrings, then there exist finitely many pairwise non-isomorphic finite simple left $R$-modules $S_{1},\ldots,S_{n}$ such that every simple left $R$-module is isomorphic to one of the $S_{i}$.
@@
    $\operatorname{Max}(R)$ and therefore $\operatorname{Max}(T)$ are finite. Moreover, $\operatorname{{Max}_{\ell}}(R)$ and $\operatorname{{Max}_{r}}(R)$ are finite.
@@
    $R/J(R)$ and therefore $T/J(T)$ are finite semisimple rings. In particular, $\operatorname{{Max}_{\ell}}(T)$ and $\operatorname{{Max}_{r}}(T)$ are finite.
@@
    For each maximal one-sided ideal $M$ of $R$, $R/M$ is finite (i.e., each simple left/right $R$-module is finite). In particular, for each maximal one-sided ideal $N$ of $T$, $T/N$ is finite (i.e., each simple left/right $T$-module is finite).
@@
(1)(1)It is obvious that for each maximal subring $S$ of $R$, $\mathbb{M}_{n}(S)$ is a maximal subring of $T$, see [1, Theorem 3.1(2)(2)]. Thus $R$ has only finitely many maximal subrings.
$(2)$ Note that for each maximal ideal $M$ of $R$, $A_{M}=\mathbb{M}_{n}(M)$ is a maximal ideal of $T$ and $T/A_{M}\cong\mathbb{M}_{n}(R/M)$. By the assumption that $T$ has only finitely many maximal subrings, we deduce that $T/A_{M}$ has only finitely many maximal subrings too. Hence by Theorem 3.5, we deduce that $R/M$ is finite (note, $n>1$ and therefore $T/A_{M}$ is not a division ring). The second part of $(2)$ is obvious by the first part of $(2)$, because each maximal ideal of $T$ is of the form $A_{M}$, for some maximal ideal $M$ of $R$.
$(3)$ With the notation of $(2)$, since $n>1$, by Theorem 2.7(4)(4), we conclude that $\mathbb{M}_{n}(R/M)$ has a maximal subring. Hence, $T$ has a maximal subring which contains $A_{M}$, say $T_{M}$. In particular, if $M\neq N\in\operatorname{Max}(R)$, then $T_{M}\neq T_{N}$, because $A_{M}+A_{N}=T$. Thus $\operatorname{Max}(R)$ is finite, a fortiori $\operatorname{Max}(T)$ is finite. For the second part of $(3)$, note that by $(1)$, $R$ has only finitely many maximal subrings, hence Proposition 3.10(1)(1), implies that $\operatorname{{Max}_{\ell}}(R)\setminus\operatorname{Max}(R)$ and $\operatorname{{Max}_{r}}(R)\setminus\operatorname{Max}(R)$ are finite. Therefore $\operatorname{{Max}_{\ell}}(R)$ and $\operatorname{{Max}_{r}}(R)$ are finite, because $\operatorname{Max}(R)$ is finite by the first part of $(3)$.
$(4)$ Since $\operatorname{{Max}_{\ell}}(R)$ is finite, we infer that $R/J(R)$ is a semisimple ring. Nothing that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring, consequently $R/J(R)$ is a finite ring. Now from the fact that $J(T)=\mathbb{M}_{n}(J(R))$, we deduce that $T/J(T)$ is finite semisimple too. Therefore $\operatorname{{Max}_{\ell}}(T)$ and $\operatorname{{Max}_{r}}(T)$ are finite.
$(5)$ It is obvious by $(4)$. ∎
We now state several corollaries.
@@
Let $R$ be a zero-dimensional ring with $\operatorname{Char}(R)=0$. Then $\mathbb{M}_{n}(R)$, where $n>1$, and $R\times R$ have infinitely many maximal subrings.
@@
Assume that $\mathbb{M}_{n}(R)$ or $R\times R$ has only finitely many maximal subrings. Then by Propositions 5.4 or 5.5, we deduce that for each maximal ideal $M$ of $R$, $R/M$ is a finite ring and therefore $\operatorname{Char}(R/M)\neq 0$. Now note that there exists a prime ideal $P$ of $R$ such that $P\cap\mathbb{Z}=0$. Since $R$ is a zero-dimensional ring we deduce that $P$ is a maximal ideal of $R$ and therefore $R/P$ is finite. But from $\mathbb{Z}\cap P=0$, we deduce that $R/P$ contains a copy of $\mathbb{Z}$, which is absurd. ∎
@@
Let $T$ be a semilocal ring (in particular, if $T$ is a left/right Artinian ring) with only finitely many maximal subrings. Then $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$, $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$, for some $n,k\geq 0$. Moreover, whenever $D_{i}$ is infinite, then $D_{i}\ncong D_{j}$ as rings, for each $j\neq i$.
@@
Since $T/J(T)$ is a semisimple ring, we deduce that $T/J(T)\cong\prod_{i\in I}\mathbb{M}_{n_{i}}(D_{i})$, where $I$ is finite, $D_{i}$ is a division ring and $n_{i}\in\mathbb{N}$. Clearly $T$ has only finitely many maximal ideals. Also note that since $T$ has only finitely many maximal subrings, then by Proposition 3.10(1)(1), we conclude that $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite. Therefore $\operatorname{{Max}_{\ell}}(T)$ is finite. Similarly $\operatorname{{Max}_{r}}(R)$ is finite. Obviously for each $i\in I$, the ring $\mathbb{M}_{n_{i}}(D_{i})$ has only finitely many maximal subrings, thus by Corollary 2.4, we deduce that whenever $i\in I$ and $n_{i}>1$, then $D_{i}$ is finite, i.e., $D_{i}$ is a finite field. Consequently $T/J(T)\cong D_{1}\times\cdots\times D_{n}\times\mathbb{M}_{n_{1}}(E_{1})\times\cdots\times\mathbb{M}_{n_{k}}(E_{k})$, where each $D_{i}$ is a division ring, $1\leq i\leq n$ and $E_{j}$ is a finite field, $1\leq j\leq k$ and $n_{j}>1$, for some $n,k\geq 0$. Finally if for some $1\leq i\neq j\leq n$, $D_{i}\cong D_{j}$ as rings, then $T$ has an ideal $I$ such that $T/I\cong D_{i}\times D_{j}\cong D_{i}\times D_{i}$. Since $T$ has only finitely many maximal subrings, we deduce that $T/I$ has only finitely many maximal subrings. Thus by Proposition 3.6, $D_{i}$ is finite, which is impossible. Hence we are done. ∎
We conclude this paper by the following immediate corollary.